Light starts in water, where the index of refraction is 1.33, and is incident on a plane interface with air, where the index is 1.00. For each of the following angles of incidence, determine the angle of refraction:
a. 35o
b. 65o
Answer #1

Refractive index of water is 1 ; n1 =1

From Snell's law, \frac{\sin i}{\sin r} = \frac{n_1}{n_2} = \frac{1}{n} ​​​​​​

Angle of Refraction, r = \sin^{-1}\left (n\sin i \right ) ​​​​​​

a.

r = \sin^{-1}\left (1.33\sin 35^o ) =49.7^o \rightarrow Answerb.

First finding the critical angle for water - air interface

C = \sin^{-1}\left (\frac{1}{n} \right) =\sin^{-1}\left (\frac{1}{1.33} \right) = 48.75^oIn the question, 65 degree is more than the critical angle, so the light beam reflects towards inside the water medium. Refraction doesn't happen here